OpenStudy (anonymous):

2 + 2+1 3/10+1 3/10=?

OpenStudy (anonymous):

1 3/10 = 13/10 2 + 2+1 3/10+1 3/10= 4+ 26/10 = 40/10 + 26/10 = 66/10 = 6 6/10